OpenStudy (anonymous):

The ratio of the height of a benzai ficus tree to the hieght of a full size ficus treen1:9 The bonsai ficusis 6 inchs tall what is the height of a full size ficus

OpenStudy (anonymous):

ratio of a:b can be written as a/b. Now 1:9 = 1/9 the question says the bonsai is of 6 inch let the full size be x Now the fraction 1/9 can be said to be equal to 6/x 1/9 = 6/x try to solve this